OpenStudy (anonymous):

can someone explain to me why cot(sin-1(1/5)=sq6/12

OpenStudy (anonymous):

\[\cot(\sin^{-1} (1/5)=\sqrt{6}/12\]

OpenStudy (anonymous):

@amistre64 @ganeshie8 @campbell_st

OpenStudy (campbell_st):

good question I drew a triangle